OpenStudy (anonymous):

Which equations are equations for the line shown? Choose all answers that are correct. A. y+3= -2(x-2) B. y+2=2(x-5) C. y-5=-2(x+2) D. y-2=-2(x+3)

OpenStudy (gabebae):

All of these are written out in point slope form. Do you know how to covert to slope intercept?

OpenStudy (anonymous):

im confused

OpenStudy (gabebae):

Point slope form is y-y1= m(x-x1) Slope intercept form is y= mx+b

OpenStudy (anonymous):

so C

OpenStudy (gabebae):

I wasn't giving you the answer I'm trying to help you get the answer

OpenStudy (gabebae):

It would be easier to graph and try out the answers by converting them to slope intercept form y=mx+b where m equals the slope and b equals the y-intercept

OpenStudy (anonymous):

Thats what I think the answer is though

OpenStudy (gabebae):

Do you know how you got the answer or did you just guess?

OpenStudy (anonymous):

my brother was explaining it to me so I did what I remember he told me to but I was just making sure

OpenStudy (gabebae):

I don't know you might need to give me a couple of minutes. I need to go afk real quick..

OpenStudy (anonymous):

ok...

OpenStudy (tkhunny):

Learn to read the point-slope form A. y+3= -2(x-2) - Passes through (2,-3). You tell me how I know that. B. y+2=2(x-5) - Passes through (5,-2). You tell me how I know that. C. y-5=-2(x+2) - Passes through (-2,5). You tell me how I know that. D. y-2=-2(x+3) - Passes through (-3,2). You tell me how I know that.

OpenStudy (anonymous):

Wait nevermind I think I got it

OpenStudy (anonymous):

I was right it was C and A